OpenStudy (anonymous):

help w/ quadratic equation: -0.01x^2+50x-62100=0

OpenStudy (anonymous):

since there is a +or- you will get 2 answers for x

OpenStudy (anonymous):

i know but i need help solving it

OpenStudy (anonymous):

-0.01x^2+50x-62100=0 x^2-5000x+6210000=0 \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Zed can help you with that, but it's literally just plugging in the values for a b c into the equation

OpenStudy (anonymous):

Here a=1, b=-5000 and c=6210000 \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]\[=\frac{5000 \pm \sqrt{(-5000)^2-4*1*6210000}}{2*1}\]\[=\frac{5000 \pm \sqrt{25000000-24840000}}{2}\]\[=\frac{5000 \pm \sqrt{160000}}{2}\]\[=\frac{5000 \pm 400}{2}\]\[=2500 \pm 200\] So x=2700 or x=2300